# thanks: ORCID ID: 0009-0003-9679-035Xthanks: ORCID ID: $0$-0001-7989-9271

### v1

@@ Abstract
We consider a simple thermodynamically consistent model that captures the interplay between autocatalytically reacting surfactants, the Marangoni effect and wetting dynamics. An ambient bath of surfactant acts as a chemostat and provides the system with chemical fuel, thereby driving it away from thermodynamic equilibrium. We find that a positive feedback loop between the local reactions and the Marangoni effect induces surface tension gradients that allow for self-propelled droplets. Besides simple directional motion, we find crawling and shuttling droplets as well as droplets performing random walks, thus exploring the entire substrate. We study the occurring dynamics and show how the observed states emerge from local and global bifurcations. Due to the underlying generic thermodynamic structure, we expect that our results are relevant not only to directly related biomimetic droplet systems but also to structurally similar systems like chemically active phase separating mixtures.
@@ 1 Introduction
The interface between hydrodynamics and chemistry is rich in fascinating phenomena, ranging from chemical gardens [1] and chemically driven active colloids [2] to periodically erupting droplets [3]. Considering this ubiquitous complexity, it seems plausible that also many biological systems are found at the intersection of these fields [4, 5, 6, 7]. Correspondingly, the study of physical processes in living matter, including hydrodynamic phenomena, has gained considerable traction in recent years. Examples include hydrodynamic models of the actomyosin complex [8, 9, 10], mitotic spindle positioning [11], chromatin dynamics [12] and of osmotic biofilm spreading [13, 14] as well as studies of protein dynamics at biomembranes [15, 16, 17, 18] and biomolecular condensates [19, 20, 21, 22, 23]. Here, minimalistic models complement complicated biologically faithful descriptions, as they are more accessible to theoretical and conceptual study. In particular, motility and self-propulsion phenomena are studied as hallmarks of active matter, with diverse underlying physico-chemical mechanisms including asymmetric distributions of catalytic activity [24, 25, 2, 26, 27, 28], self-induced wettability gradients [29, 30, 31, 32, 33, 34], active stresses [35, 36, 37, 38] and enzymatically maintained concentration gradients [22, 23]. Among such self-propulsion strategies, the (solutal) Marangoni effect, i.e., forces localized at interfaces that result from surface tension gradients, is especially common in the context of biomimetic and prebiotic systems. It is employed in, e.g., drop-based microswimmers [39, 40] and some simple models of protocell motility [41]. Aimed at achieving biomimetic functions, previous studies have combined autocatalytic pattern-forming reactions like the Belousov-Zhabotinsky reaction with (droplet) hydrodynamics [42, 43, 44, 45] resulting in complex interfacial deformation and motility modes due to chemically driven Marangoni flows. Autocatalytic mechanisms are also discussed as possible forms of molecular self-replication under prebiotic conditions [46, 47, 48]. Interestingly, candidates for autocatalytic self-replication also include, e.g., amphiphilic peptides [49] that can adsorb at water-air interfaces. In view of these observations, it is compelling to study whether the coupling of autocatalytic processes at interfaces and droplet hydrodynamics can result in complex forms of motility without the highly specialized biochemical machinery of biological cells.
@@
The present article is organized as follows. In Sec. 2 we provide a thermodynamically consistent free-energy based description of droplets covered by chemically reacting surfactants. We discuss how the presence of chemical fuel results in persistent nonequilibrium behavior. In Sec. 3 we first investigate droplets on a one-dimensional substrate and study the underlying self-propulsion mechanism. In Sec. 4 we study more complex forms of droplet motion, namely periodic stick-slip-like motion (“crawling”) and back-and-forth motion (“shuttling”), and the related bifurcations. Finally, we briefly turn to droplets on a two-dimensional substrate, where the additional degree of freedom allows for highly complex types of motility. In Sec. 5, we recapitulate our findings, discuss potential implications for biomimetic and related biological or prebiotic systems and list possible extensions of the model.
@@ 2 Droplets covered by autocatalytic surfactants
where $\mathcal{S}$ denotes the substrate plane, $\xi=\sqrt{1+||\nabla h||^{2}}$ is the local metric factor of the droplet surface with $\nabla=(\partial_{x},\partial_{y})^{\text{T}}$ and |||||$\cdot$||| | ⋅ | | is the Euclidean norm. The partial derivative with respect to $i$ is denoted by $\partial_{i}$. Equation (1) comprises two contributions, namely the mesocopic wetting energy $f(h)$ that encodes all liquid-substrate interactions and the surface energy $g(\Gamma_{1},\Gamma_{2})$. We here neglect possible surfactant-substrate interactions [50], such that $f$ is independent of $\Gamma_{1}$ and $\Gamma_{2}$. Conversely, $g$ is assumed to be independent of $h$. Specifically, we assume that the wetting energy is a simple superposition of long-range attractive and short-range repulsive interactions,
@@
where $\gamma$ symmetrically depends on both species, for details see [54, 55]. This implies that only the total surfactant count determines the surface tension and therefore excludes (self-propulsion) effects based on differences in surfactant properties [39, 40]. The surfactants chemically react in the reversible autocatalytic reaction
@@
where $\Gamma_{1}$ catalyzes its own production with the reaction rate $r>0$. Consequently, we refer to $\Gamma_{1}$ as the autocatalyst and to $\Gamma_{2}$ as the reactant. The droplet also exchanges surfactant with the ambient bath which acts as a reservoir or chemostat for both chemical species. This is modeled as reversible adsorption-desorption reactions,
@@
where forward reactions (e.g., $\Gamma_{1}\rightarrow\text{bath}$) correspond to desorption from the free surface and $\beta_{1},\beta_{2}>0$ are the reaction rates for $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Instead of treating an ambient bath, one could equivalently assume an exchange with the droplet bulk as a reservoir. We assume that the reservoir is large such that the exchange of surfactant with the droplet does not significantly alter the concentrations of species in the bath and the corresponding constant chemical potentials $\mu_{1}$ and $\mu_{2}$ are convenient control parameters. The reactions (6) can therefore continuously provide the droplet with chemical fuel and drive the system away from thermodynamic equilibrium. The dynamics is then described by a “passive core” in gradient dynamics form with thermodynamic forces derived from variations of the free energy (1), that is augmented by the exchange reactions with the bath. The complete model reads
$\partial_{t}h$ $=-\nabla\cdot\bm{j}_{h}$ $=\nabla\cdot\left[Q_{hh}\nabla\frac{\delta F}{\delta h}+Q_{h\Gamma_{1}}\nabla\frac{\delta F}{\delta\tilde{\Gamma}_{1}}+Q_{h\Gamma_{2}}\nabla\frac{\delta F}{\delta\tilde{\Gamma}_{2}}\right],$
$\partial_{t}\tilde{\Gamma}_{1}$ $=-\nabla\cdot\bm{j}_{1}+\mathcal{R}+\mathcal{B}_{1}$ $=\nabla\cdot\left[Q_{\Gamma_{1}h}\nabla\frac{\delta F}{\delta h}+Q_{\Gamma_{1}\Gamma_{1}}\nabla\frac{\delta F}{\delta\tilde{\Gamma}_{1}}+Q_{\Gamma_{1}\Gamma_{2}}\nabla\frac{\delta F}{\delta\tilde{\Gamma}_{2}}\right]+\mathcal{R}+\mathcal{B}_{1},$ (7)
$\partial_{t}\tilde{\Gamma}_{2}$ $=-\nabla\cdot\bm{j}_{2}-\mathcal{R}+\mathcal{B}_{2}$ $=\nabla\cdot\left[Q_{\Gamma_{2}h}\nabla\frac{\delta F}{\delta h}+Q_{\Gamma_{2}\Gamma_{1}}\nabla\frac{\delta F}{\delta\tilde{\Gamma}_{1}}+Q_{\Gamma_{2}\Gamma_{2}}\nabla\frac{\delta F}{\delta\tilde{\Gamma}_{2}}\right]-\mathcal{R}+\mathcal{B}_{2}.$
Here, $\tilde{\Gamma}_{1,2}=\xi\Gamma_{1,2}$ are “projected” densities (particles per unit substrate area) that directly correspond to the particle numbers of surfactant and are independent of the surface geometry, for details see [54, 56, 55]. For a general account of the thermodynamic structure of Eqs. (7) we also refer to [55] and references therein. The fully equivalent hydrodynamic form is given in Appendix A. Note that Eqs. (7) conserve the total liquid volume. The variational derivatives $\delta F/\delta h$ and $\delta F/\delta\tilde{\Gamma}_{1},\delta F/\delta\tilde{\Gamma}_{2}$ correspond to the liquid pressure and the chemical potentials of $\Gamma_{1}$ and $\Gamma_{2}$ on the droplet surface, respectively. The transport fluxes $\bm{j}_{h}$ and $\bm{j}_{1},\bm{j}_{2}$ in Eqs. (7) are thus linear in gradients of pressure and chemical potentials and represent diffusive and advective contributions to transport, including Marangoni fluxes. The associated mobility matrix [54, 55],
@@
is positive definite and symmetric, ensuring non-negative entropy production and satisfying the Onsager relations [57, 58, 59]. It corresponds to a thin-film description for droplets covered by insoluble surfactants without slip at the substrate [54, 55]. Here, $\eta>0$ is the dynamic viscosity of the liquid and $D_{1},D_{2}>0$ are diffusive mobilities of the respective surfactant. Note that in the underlying hydrodynamic problem, the transport fluxes $\bm{j}_{h}$ and $\bm{j}_{1},\bm{j}_{2}$ respectively correspond to the vertically intergrated horizontal components of the liquid velocity and to diffusive and advective surfactant transport at the free surface [60]. Equation (8) implies that the ambient medium does not contribute to the drop dynamics, e.g., because it is of low viscosity. This differs from the commonly treated Marangoni-driven microswimmers [39, 40]. However, this assumption could be relaxed in the future by incorporating aspects of [61].
@@
$\mathcal{R}=r\left[\exp\left(\frac{2}{k_{b}T}\frac{\delta F}{\delta\tilde{\Gamma}_{1}}+\frac{1}{k_{b}T}\frac{\delta F}{\delta\tilde{\Gamma}_{2}}\right)-\exp\left(\frac{3}{k_{b}T}\frac{\delta F}{\delta\tilde{\Gamma}_{1}}\right)\right],$ (9)
@@
$\mathcal{B}_{1}$ $=\beta_{1}\left[\exp\left(\frac{\mu_{1}}{k_{b}T}\right)-\exp\left(\frac{1}{k_{b}T}\frac{\delta F}{\delta\tilde{\Gamma}_{1}}\right)\right],$ (10)
$\mathcal{B}_{2}$ $=\beta_{2}\left[\exp\left(\frac{\mu_{2}}{k_{b}T}\right)-\exp\left(\frac{1}{k_{b}T}\frac{\delta F}{\delta\tilde{\Gamma}_{2}}\right)\right],$
where $\mu_{1},\mu_{2}$ are the uniform and constant chemical potentials of $\Gamma_{1}$ and $\Gamma_{2}$ in the bath that acts as a chemostat. This reflects that the concentrations in the bath are not significantly affected by the exchange with the droplet. Note that in Sec. 4.2 of Ref. [55], the linearized expressions of Eqs. (10) are employed, which are strictly valid only near thermodynamic equilibrium [57]. For ${\mu_{1}=\mu_{2}}$, Eqs. (7) represent an open albeit passive system where the grand potential $F-\int\sum_{i}\xi\mu_{i}\Gamma_{i}\text{d}x^{2}$ continuously decreases. [^1] See, e.g., the formulation of the second law of thermodynamics for open reaction-diffusion systems in Ref. [64] which also applies here. Then, the system ultimately relaxes to a single droplet that is uniformly covered by both surfactants. Otherwise, $\mu_{1}$ and $\mu_{2}$ reflect driving forces and the system is permanently out-of-equilibrium. More generally, the nonequilibrium driving results from an incompatibility between the chemical potentials of the chemostat with respect to the conditions for thermodynamic equilibrium [64]. Thus, if the chemical potentials of the chemostat are uniform and constant, two or more chemostatted species are necessary for sustained nonequilibrium. Therefore, systems with only one chemostat such as the chemically driven running drop discussed in Sec. 4.1 of Ref. [55] indeed relax toward thermodynamic equilibrium, i.e. the (semi-)grand potential continuously decreases [64] (in contrast to the increasing free energy shown in Ref. [55]). However, if the running drop moves in an infinite domain the actual equilibrium is never reached somewhat similarly to a drop sliding down an infinite incline.
To study the emerging dynamics, we nondimensionalize Eqs. (7). Importantly, we explicitly use the assumption that slopes in the droplet profile are small ($||\nabla h||\ll 1$), resulting in the approximation $\tilde{\Gamma}_{1,2}\approx\Gamma_{1,2}$ in the final equations. For details, we refer to Appendix B and note that from hereon all quantities are nondimensional. In the following, we investigate the dynamics of self-propelled droplets which occur for sufficiently strong nonequilibrium driving.

### v3

# thanks: ORCID ID: 0009-0003-9679-035Xthanks: ORCID ID: 0000-0001-7989-9271

@@ Abstract
We consider a simple thermodynamically consistent model that captures the self-organized chemomechanical coupling resulting from the interplay between autocatalytically reacting surfactants, the Marangoni effect and wetting dynamics. An ambient bath of surfactant acts as a chemostat and provides the system with chemical fuel, thereby driving it away from thermodynamic equilibrium. We find that a positive feedback loop between the local reactions and the Marangoni effect induces surface tension gradients that allow for self-propelled droplets. Besides simple directional motion, we find crawling and shuttling droplets as well as droplets performing random walks, thus exploring the entire substrate. We study the occurring chemomechanical motility modes and show how the observed dynamic states emerge from local and global bifurcations. Due to the underlying generic thermodynamic structure, we expect that our results are relevant not only to directly related biomimetic droplet systems but also to structurally similar systems like chemically active phase-separating mixtures.
@@ 1 Introduction
The interface between hydrodynamics and chemistry is rich in fascinating phenomena, ranging from chemical gardens [1] and chemically driven active colloids [2] to periodically erupting droplets [3]. Considering this ubiquitous complexity, it seems plausible that also many biological systems are found at the intersection of these fields [4, 5, 6, 7]. Correspondingly, the study of physical processes in living matter, including hydrodynamic phenomena, has gained considerable traction in recent years. Examples include hydrodynamic models of the actomyosin complex [8, 9, 10], mitotic spindle positioning [11], chromatin dynamics [12], protoplasmic droplets [13, 14] and osmotic biofilm spreading [15, 16] as well as studies of reaction- and diffusion-based protein dynamics at biomembranes [17, 18, 19, 20] and biomolecular condensates [21, 22, 23, 24, 25]. Here, minimalistic models complement complicated biologically faithful descriptions, as they are more accessible to theoretical and conceptual study. In particular, motility and self-propulsion phenomena are studied as hallmarks of active matter, with diverse underlying physico-chemical mechanisms including asymmetric distributions of catalytic activity [26, 27, 2, 28, 29, 30], self-induced wettability gradients [31, 32, 33, 34, 35, 36], active stresses [13, 14, 37, 38, 39, 40] and enzymatically maintained concentration gradients [24, 25]. Among such self-propulsion strategies, the (solutal) Marangoni effect, i.e., the emergence of mechanical forces localized at interfaces that results from surface tension gradients, is especially common in the context of biomimetic and prebiotic systems. It is employed in, e.g., drop-based microswimmers [41, 42] and some simple models of protocell motility [43]. Aimed at achieving biomimetic functions, previous studies have combined autocatalytic pattern-forming reactions like the Belousov-Zhabotinsky reaction with (droplet) hydrodynamics [44, 45, 46, 47] and polymeric gels [48, 49, 50] resulting in complex deformation and motility modes due to chemically driven flows. Autocatalytic mechanisms are also discussed as possible forms of molecular self-replication under prebiotic conditions [51, 52, 53]. Interestingly, candidates for autocatalytic self-replication also include, e.g., amphiphilic peptides [54] that can adsorb at water-air interfaces. In view of these observations, it is compelling to study whether the coupling of autocatalytic processes at interfaces and droplet hydrodynamics can result in complex forms of motility without the highly specialized biochemical machinery of biological cells.
@@
We remark that the model briefly considered in Sec. 4.2 of Ref. [55] by the authors is related to the one presented in this article. A notable difference consists in the modeling of the exchange reactions with the chemostats (which provide the chemical fuel) which are here, in contrast to Ref. [55], not linearized about chemical equilibrium. More importantly, here, we consider parameter regimes for which the reaction-diffusion subsystem of surfactants without any coupling to hydrodynamics does not form any patterns. Instead, the various modes of droplet motion emerge only as a consequence of a chemomechanical feedback loop that results from the interplay of nonlinear chemical reactions and the Marangoni effect which we describe in Sec. 3.
The present article is organized as follows. In Sec. 2 we provide a thermodynamically consistent free-energy based description of droplets covered by chemically reacting surfactants. We discuss how the presence of chemical fuel results in persistent nonequilibrium behavior. In Sec. 3 we first investigate droplets on a one-dimensional substrate and study the underlying self-propulsion mechanism which we identify as a positive feedback loop between the local chemical reactions and the Marangoni effect. In Sec. 4 we study more complex forms of droplet motion, namely periodic stick-slip-like motion (“crawling”) and back-and-forth motion (“shuttling”), and the related bifurcations. Finally, we briefly turn to droplets on a two-dimensional substrate, where the additional degree of freedom allows for highly complex types of motility. In Sec. 5, we recapitulate our findings, discuss potential implications for biomimetic and related biological or prebiotic systems and list possible extensions of the model.
@@ 2 Droplets covered by autocatalytic surfactants
[…]
@@
where the surface tension $\gamma$ symmetrically depends on both species, for details see [60, 55]. This implies that only the total surfactant count determines the surface tension and therefore excludes (self-propulsion) effects based on differences in surfactant properties [41, 42]. The surfactants chemically react in the reversible autocatalytic reaction
@@
where $\Gamma_{1}$ catalyzes its own production with the reaction rate $r>0$. Consequently, we refer to $\Gamma_{1}$ as the autocatalyst and to $\Gamma_{2}$ as the reactant. Reaction (5) corresponds to the central nonlinear reactions of paradigmatic pattern forming systems like the Brusselator [61, 62] and the Schlögl model [62]. Here, it represents a simple incorporation of a chemical nonlinearity. In systems containing multiple surfactants, such higher order reactions may be associated with the formation of micelles that consist of both species [63, 64]. The droplet also exchanges surfactant with the ambient bath which acts as a reservoir or chemostat for both chemical species. This is modeled as reversible adsorption-desorption reactions,
@@
where forward reactions (e.g., $\Gamma_{1}\rightarrow\text{bath}$) correspond to desorption from the free surface and $\beta_{1},\beta_{2}>0$ are the reaction rates for $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Instead of treating an ambient bath, one could equivalently assume an exchange with the droplet bulk as a reservoir. In both cases, we assume that the reservoir is large such that the exchange of surfactant with the drop surface does not significantly alter the concentrations of species in the bath and the corresponding constant chemical potentials $\mu_{1}$ and $\mu_{2}$ are convenient control parameters. For a thermodynamically sound justification of this approximation, see Ref. [65]. The reactions (6) can therefore continuously provide the droplet with chemical fuel and drive the system away from thermodynamic equilibrium. The dynamics is then described by a “passive core” in gradient dynamics form with thermodynamic forces derived from variations of the free energy (1), that is augmented by the exchange reactions with the bath. The complete model reads
$\partial_{t}h$ $=-\nabla\cdot\bm{j}_{h}$ $=\nabla\cdot\left[Q_{hh}\nabla\frac{\delta F}{\delta h}+Q_{h\Gamma_{1}}\nabla\frac{\delta F}{\delta\hat{\Gamma}_{1}}+Q_{h\Gamma_{2}}\nabla\frac{\delta F}{\delta\hat{\Gamma}_{2}}\right],$
$\partial_{t}\hat{\Gamma}_{1}$ $=-\nabla\cdot\bm{j}_{1}+\mathcal{R}+\mathcal{B}_{1}$ $=\nabla\cdot\left[Q_{\Gamma_{1}h}\nabla\frac{\delta F}{\delta h}+Q_{\Gamma_{1}\Gamma_{1}}\nabla\frac{\delta F}{\delta\hat{\Gamma}_{1}}+Q_{\Gamma_{1}\Gamma_{2}}\nabla\frac{\delta F}{\delta\hat{\Gamma}_{2}}\right]+\mathcal{R}+\mathcal{B}_{1},$ (7)
$\partial_{t}\hat{\Gamma}_{2}$ $=-\nabla\cdot\bm{j}_{2}-\mathcal{R}+\mathcal{B}_{2}$ $=\nabla\cdot\left[Q_{\Gamma_{2}h}\nabla\frac{\delta F}{\delta h}+Q_{\Gamma_{2}\Gamma_{1}}\nabla\frac{\delta F}{\delta\hat{\Gamma}_{1}}+Q_{\Gamma_{2}\Gamma_{2}}\nabla\frac{\delta F}{\delta\hat{\Gamma}_{2}}\right]-\mathcal{R}+\mathcal{B}_{2}.$
Here, $\hat{\Gamma}_{1,2}=\xi\Gamma_{1,2}$ are “projected” densities (particles per unit substrate area) that directly correspond to the particle numbers of surfactant and are independent of the surface geometry, for details see [60, 66, 55]. For a general account of the thermodynamic structure of Eqs. (7) we also refer to [55] and references therein. The fully equivalent hydrodynamic form is given in Appendix A. Note that Eqs. (7) conserve the total liquid volume. The variational derivatives $\delta F/\delta h$ and $\delta F/\delta\hat{\Gamma}_{1},\delta F/\delta hat{\Gamma}_{2}$ correspond to the liquid pressure and the chemical potentials of $\Gamma_{1}$ and $\Gamma_{2}$ on the droplet surface, respectively. The transport fluxes $\bm{j}_{h}$ and $\bm{j}_{1},\bm{j}_{2}$ in Eqs. (7) are thus linear in gradients of pressure and chemical potentials and represent diffusive and advective contributions to transport, including Marangoni fluxes. The associated mobility matrix [60, 55],
@@
is positive definite and symmetric, ensuring non-negative entropy production and satisfying the Onsager relations [67, 68, 69]. It corresponds to a thin-film description for droplets covered by insoluble surfactants without slip at the substrate [60, 55]. Here, $\eta>0$ is the dynamic viscosity of the liquid and $D_{1},D_{2}>0$ are diffusive mobilities of the respective surfactant. Note that in the underlying hydrodynamic problem, the transport fluxes $\bm{j}_{h}$ and $\bm{j}_{1},\bm{j}_{2}$ respectively correspond to the vertically integrated horizontal components of the liquid velocity and to diffusive and advective surfactant transport at the free surface [70]. Equation (8) implies that the ambient medium does not contribute to the drop dynamics, e.g., because it is of low viscosity. This differs from the commonly treated Marangoni-driven microswimmers [41, 42]. However, this assumption could be relaxed in the future by incorporating aspects of [71]. The model (7) could also be extended to treat soluble surfactants [66], e.g., to capture a possible depletion or enrichment of surfactants in the bath. Then, the chemical potentials $\mu_{1},\mu_{2}$ would be dynamical quantities, instead of the explicit chemostatting assumed here.
@@
$\mathcal{R}=r\left[\exp\left(\frac{2}{k_{b}T}\frac{\delta F}{\delta\hat{\Gamma}_{1}}+\frac{1}{k_{b}T}\frac{\delta F}{\delta\hat{\Gamma}_{2}}\right)-\exp\left(\frac{3}{k_{b}T}\frac{\delta F}{\delta\hat{\Gamma}_{1}}\right)\right],$ (9)
@@
$\mathcal{B}_{1}$ $=\beta_{1}\left[\exp\left(\frac{\mu_{1}}{k_{b}T}\right)-\exp\left(\frac{1}{k_{b}T}\frac{\delta F}{\delta\hat{\Gamma}_{1}}\right)\right],$ (10)
$\mathcal{B}_{2}$ $=\beta_{2}\left[\exp\left(\frac{\mu_{2}}{k_{b}T}\right)-\exp\left(\frac{1}{k_{b}T}\frac{\delta F}{\delta\hat{\Gamma}_{2}}\right)\right],$
where $\mu_{1},\mu_{2}$ are the uniform and constant chemical potentials of $\Gamma_{1}$ and $\Gamma_{2}$ in the bath that acts as a chemostat. This reflects that the concentrations in the bath are not significantly affected by the exchange with the drop surface. Note that in Sec. 4.2 of Ref. [55], the linearized expressions of Eqs. (10) are employed, which are strictly valid only near thermodynamic equilibrium [67]. For $\mu_{1}=\mu_{2}$}, Eqs. (7) represent an open albeit passive system where the grand potential $F-\int\sum_{i}\xi\mu_{i}\Gamma_{i}\text{d}^{2}x$ continuously decreases. [^1] See, e.g., the formulation of the second law of thermodynamics for open reaction-diffusion systems in Ref. [74] which also applies here. Then, the system ultimately relaxes to a single droplet that is uniformly covered by both surfactants. Otherwise, $\mu_{1}$ and $\mu_{2}$ reflect driving forces and the system is permanently out-of-equilibrium. More generally, the nonequilibrium driving results from an incompatibility between the chemical potentials of the chemostat with respect to the conditions for thermodynamic equilibrium [74]. Thus, if the chemical potentials of the chemostat are uniform and constant, two or more chemostatted species are necessary for sustained nonequilibrium. Therefore, systems with only one chemostat such as the chemically driven running drop discussed in Sec. 4.1 of Ref. [55] indeed relax toward thermodynamic equilibrium, i.e. the (semi-)grand potential continuously decreases [74] (in contrast to the increasing free energy shown in Ref. [55]). However, if the running drop moves in an infinite domain the actual equilibrium is never reached somewhat similar to a drop sliding down an infinite incline.
To study the emerging dynamics, we nondimensionalize Eqs. (7). Importantly, we explicitly use the assumption that slopes in the droplet profile are small ($\nabla h||\ll 1$), resulting in the approximation $\hat{\Gamma}_{1,2}\approx\Gamma_{1,2}$ in the final equations. For details, we refer to Appendix B, where we also discuss possible dimensional parameter values. From hereon, all quantities are nondimensional. In the following, we investigate the dynamics of self-propelled droplets which occur for sufficiently strong nonequilibrium driving.